OpenStudy (anonymous):

You roll two dice. What is the probability of rolling a 3 or a 5 on the second die, given that you rolled an odd number on the first die? A 6X6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

2 out of six or 2/6. The first roll has no effect on the outcome of the second.

OpenStudy (anonymous):

ok awesome thanks!

OpenStudy (anonymous):

2/6 alwats since the events are independent

OpenStudy (amistre64):

out of 36 outcomes; we get 6*1 chances out of 36 to get a 3 6*1 chances out of 36 to get a 5 6+6 = 12 chances out of 36 to get a 3 or 5 on the 2nd die 12/36 = 2/6 = 1/3 if i did it right